# Merlin: Multi-Stage Curriculum Alignment for Multilingual Encoder and LLM Fusion

### v2

@@ 5 Analysis
From layers 18–28 MERLIN reaches its peak sentence–level retrieval@5 (~0.30). The rise coincides with the PEFT fine–tuning in Stage II, which adjusts a small set of low-rank weights on top of the frozen LLM while keeping the connector $f_{\sigma^{\star}}$ fixed. This adaptation strengthens the multilingual abstractions that naturally emerge in the decoder’s middle blocks—the region reported to be the most language-agnostic once parallel supervision is available (Liu and Niehues, 2025). As a result, non-English queries are already embedded in the English reasoning sub-space before the later, generation-oriented layers are reached, providing a solid cross-lingual scaffold for downstream mathematical reasoning tasks such as AfriMGSM.
@@ 6 Related Work
LRLs often suffer from limited annotated data and high-quality machine translation resources, creating significant challenges for NLP models (Deshpande et al., 2024; Qin et al., 2024; Tonja et al., 2024; Khan et al., 2023; Magueresse et al., 2020). Moreover, these languages can exhibit complex morphological structures and orthographic variations (Ghosh et al., 2024; Nzeyimana, 2024; Lopo and Tanone, 2024; Issaka et al., 2024; Lusito et al., 2023), posing further challenges. While multilingual models can handle many languages, they frequently underperform in truly low-resource settings and may even trail behind strong monolingual baselines (Yoon et al., 2024; Du et al., 2024; Blevins et al., 2024). Existing efforts to boost LRL performance focus primarily on dataset creation (Adelani et al., 2021; Muhammad et al., 2022; Adelani et al., 2024a; Yong et al., 2024) or adaptation of pretrained models (Alabi et al., 2022; Wu et al., 2024; Csaki et al., 2023), yet these strategies remain constrained by data scarcity and fail to fully address complex linguistic properties (Urbizu et al., 2023; Wang et al., 2023).
@@
Recent efforts to enhance multilingual reasoning in LLMs include carefully devised prompt engineering to facilitate multi-step inference or chain-of-thought reasoning (Huang et al., 2023; Qin et al., 2023), although these techniques often lose effectiveness with smaller open-source models (Touvron et al., 2023). Another approach involves supervised fine-tuning on translated data (Chen et al., 2024; Zhu et al., 2024), which infuses language or task-specific knowledge into the model. Building upon this, Fan et al. (2025) extend the work of Zhu et al. (2024) by optimizing both the training parameters and the training stages focusing on language-specific neurons. Their improvements lead to higher average performance on multilingual mathematical reasoning benchmarks. However, reliance on machine translation can introduce errors (Shi et al., 2022) and may overlook the LLM’s built-in multilingual capabilities. To circumvent these issues, MindMerger (Huang et al., 2024) sidesteps autoregressive translation by directly integrating an external multilingual encoder, preserving the LLM’s internal parameters.
@@ 7 Conclusion
This work introduced MERLIN, a lightweight framework for strengthening multilingual mathematical reasoning and natural-language inference in large language models. A two-stage curriculum first learns a compact connector that projects multilingual encoder states into the frozen LLM embedding space, then refines the decoder through DoRA-based task specialisation. The approach preserves the original English reasoning capabilities while substantially improving performance in both mid- and low-resource languages.
Empirical evaluation on four benchmarks demonstrates that MERLIN outperforms strong open-source systems—including MindMerger, SLAM, and LangBridge—on MGSM, MSVAMP, AfriMGSM and AfriXNLI, and narrows the gap to closed models such as GPT-4o.
@@ Appendix A Dataset-Size Ablation
Table 5: MERLIN performance on MGSM and MSVAMP across synthetic-training set sizes. Bn/Th/Sw are mid-resource (MRL); the remaining seven languages are high-resource (HRL).
@@ Appendix D Ablation of Training Adapters
Table 8 contrasts two parameter-efficient adapters applied to the decoder. LoRA (Hu et al., 2022) injects trainable low-rank update matrices into the linear weights, while all original parameters stay frozen. DoRA (Liu et al., 2024) keeps the same low-rank structure but factorises each weight into a direction and a magnitude, yielding the same footprint yet a more explicit decomposition.
@@ Appendix E AmericasNLI Evaluation
To assess whether the mapping curriculum generalises beyond the African and European test beds, we evaluate MERLIN on AmericasNLI, a three-way NLI benchmark that targets Aymara, Guarani, and Quechua. For MERLIN and the MindMerger baseline, we follow the same data pipeline as for AfriXNLI: 3,000 English premise–hypothesis pairs are sampled from MultiNLI, translated into each target language with NLLB 600m distilled, and used as supervision for Stage I.b, Stage I.c, and Stage II. [^4] No task-specific data from the original AmericasNLI development set is seen during training. The non-adapted LLaMAX2-7B-XNLI (Lu et al., 2024) serves as the baseline decoder-only model. Results are summarised in Table 9.
@@ Appendix F Training Configurations
All parameter-efficient runs (MERLIN, QAlign-LoRA) employ rank-16 LoRA/DoRA adapters injected into the self-attention q_proj and v_proj projections:

### v3

@@ 5 Analysis of Language alignment and Effects on multilingualism
From layers 18–28 Merlin reaches its peak sentence–level retrieval@5 (~0.30). The rise coincides with the PEFT fine–tuning in Stage II, which adjusts a small set of low-rank weights on top of the frozen LLM while keeping the connector $f_{\sigma^{\star}}$ fixed. This adaptation strengthens the multilingual abstractions that naturally emerge in the decoder’s middle blocks—the region reported to be the most language-agnostic once parallel supervision is available (liu2025middlelayerrepresentationalignmentcrosslingual). As a result, non-English queries are already embedded in the English reasoning sub-space before the later, generation-oriented layers are reached, providing a solid cross-lingual scaffold for downstream mathematical reasoning tasks such as AfriMGSM.
@@ 6 Related Work
LRLs often suffer from limited annotated data and high-quality machine translation resources, creating significant challenges for NLP models (magueresse2020lowresourcelanguagesreviewpast; lee_pre-trained_2022; app13158566; deshpande2024chainoftranslationpromptingcotrnovel; qin2024multilinguallargelanguagemodel; tonja-etal-2024-nlp). Moreover, these languages can exhibit complex morphological structures and orthographic variations (ghosh2024morphologybasedinvestigationpositionalencodings; nzeyimana2024lowresourceneuralmachinetranslation; lopo2024constructingexpandinglowresourceunderrepresented; issaka2024ghanaiannlplandscapelook; lusito2023textnormalizationlowresourcelanguages), posing further challenges. While multilingual models can handle many languages, they frequently underperform in truly low-resource settings and may even trail behind strong monolingual baselines (yoon-etal-2024-langbridge; du2024chinesetinyllmpretraining; blevins-etal-2024-breaking). Existing efforts to boost LRL performance focus primarily on dataset creation (adelani-etal-2021-masakhaner; muhammad2022naijasentinigeriantwittersentiment; adelani-etal-2024-sib; yong2024lexcgengeneratingdataextremely) or adaptation of pretrained models (alabi-etal-2022-adapting; wu2024adaptinglargelanguagemodels; csaki2023efficientlyadaptingpretrainedlanguage; schmidt-etal-2024-self), yet these strategies remain constrained by data scarcity and fail to fully address complex linguistic properties (urbizu-etal-2023-scaling; wang2023challengingbenchmarklowresourcelearning).
@@
Recent efforts to enhance multilingual reasoning in LLMs include carefully devised prompt engineering to facilitate multi-step inference or chain-of-thought reasoning (huang-etal-2023-languages; qin-etal-2023-cross), although these techniques often lose effectiveness with smaller open-source models (touvron2023llama2openfoundation). Another approach involves supervised fine-tuning on translated data (chen2024breakinglanguagebarriersmultilingual; zhu2024questiontranslationtrainingbetter), which infuses language or task-specific knowledge into the model. Building upon this, fan2025slamefficientmultilingualreasoning extend the work of zhu2024questiontranslationtrainingbetter by optimizing both the training parameters and the training stages focusing on language-specific neurons. Their improvements lead to higher average performance on multilingual mathematical reasoning benchmarks. However, reliance on machine translation can introduce errors (shi2022languagemodelsmultilingualchainofthought) and may overlook the LLM’s built-in multilingual capabilities. To circumvent these issues, MindMerger (huang2024mindmergerefficientboostingllm) sidesteps autoregressive translation by directly integrating an external multilingual encoder, preserving the LLM’s internal parameters.
@@ 7 Conclusion
This work introduced Merlin, a lightweight framework for strengthening multilingual mathematical reasoning and natural‑language inference in large language models. A two‑stage curriculum first learns a compact connector that projects multilingual encoder states into the frozen LLM embedding space, then refines the decoder through DoRA‑based task specialization. The approach preserves the original English reasoning capabilities while substantially improving performance in both mid‑ and low‑resource languages.
Our analysis shows that model merging enhances cross-lingual alignment across layers compared to using LLMs alone. In particular, our framework, Merlin, improves alignment within the model’s internal reasoning subspace, yielding stronger performance on reasoning tasks. We also find that the choice of encoder is critical for low-resource languages: encoders with greater exposure to LRL data produce representations that align more closely with the LLM’s internal language (often English), thereby supplying richer cross-lingual signals for model merging.
Empirical evaluation on four benchmarks demonstrates that Merlin outperforms strong open‑source systems—including MindMerger, SLAM, and LangBridge—on MGSM, MSVAMP, AfriMGSM and AfriXNLI, and narrows the gap to closed models such as GPT‑4o.
@@ Appendix A Dataset-Size Ablation
Table 5: Merlin performance on MGSM and MSVAMP across synthetic‑training set sizes. Bn/Th/Sw are mid‑resource (MRL); the remaining seven languages are high‑resource (HRL).
@@ Appendix D Ablation of Training Adapters
Table 8 contrasts two parameter-efficient adapters applied to the decoder. LoRA (hu2022lora) injects trainable low-rank update matrices into the linear weights, while all original parameters stay frozen. DoRA (liu2024doraweightdecomposedlowrankadaptation) keeps the same low-rank structure but factorises each weight into a direction and a magnitude, yielding the same footprint yet a more explicit decomposition.
@@ Appendix E AmericasNLI Evaluation
To assess whether the mapping curriculum generalises beyond the African and European test beds, we evaluate Merlin on AmericasNLI, a three-way NLI benchmark that targets Aymara, Guarani, and Quechua. For Merlin and the MindMerger baseline, we follow the same data pipeline as for AfriXNLI: 3,000 English premise–hypothesis pairs are sampled from MultiNLI, translated into each target language with NLLB 600m distilled, and used as supervision for Stage I.b, Stage I.c, and Stage II. [^3] No task-specific data from the original AmericasNLI development set is seen during training. The non-adapted LLaMAX2-7B-XNLI (lu-etal-2024-llamax) serves as the baseline decoder-only model. Results are summarised in Table 9.
@@ Appendix F Training Configurations
All parameter-efficient runs (Merlin, QAlign-LoRA) employ rank-16 LoRA/DoRA adapters injected into the self‐attention q_proj and v_proj projections: